OpenStudy (ejayy12):

Please please please, Help me out! Which of the following equations is of a parabola with a vertex at (1, -2)? a.y = (x - 1)^2 - 2 b.y = (x - 1)^2 + 2 c.y = (x + 1)^2 - 2 d.y = (x + 1)^2 + 2

OpenStudy (ejayy12):

is it C?

OpenStudy (antiprime):

ok so what u wanna do is since x is 1 it will be -1

OpenStudy (ejayy12):

so B?

OpenStudy (welshfella):

compare each choice with the general forn y = (x - b)^2 + c where (b,c) is the vertex

OpenStudy (ejayy12):

-1?

OpenStudy (welshfella):

no compare the options with general form

OpenStudy (welshfella):

y = ( x - b)^2 + c b = 1 and c = -2 y = ( x - 1)^2 - 2

OpenStudy (knov):